OpenStudy (anonymous):

a ship sails 30km on a bearing of 220degrees, then 20km on a bearing of 250degrees. These are true bearings. find how far south of the original position it is

OpenStudy (anonymous):

...

OpenStudy (anonymous):

nearly done..

OpenStudy (campbell_st):

you need to draw a picture showing the information  you need to find the distance x and the distance y, which will total to the distance south of the starting point.... the bearings 220 degrees is 40 degrees west of south... so the angle to use will be 40 the bearinf 250 is 70 west of south...this will be used to find y. the respective hypotenuses are 30km and 20km find x \[\cos(40) = \frac{x}{30}...or.... x = 30 \times \cos(40)\] to find y \[\cos(70) = \frac{ y}{20}...or... y = 20 \times \cos(70)\] and x and y together for the total distance south of the starting point.